Rob is going to the country fair. There are two ticket options as shown in the chart below.
Ticket A= admissions price= $4.00, price per ride= $0.40
Ticket B= admissions price= $2.00, price per ride= $0.80
Find the number of rides for which the total cost is the same with both ticket options

Answer by ikleyn(52915)

.
A = 400 + 40*r = 200 + 80*r = B    (cost in cents)


Solve this equation

400 + 40r = 200 + 80r.


Can yo do it on your own ?


If not, then read the lesson
    - HOW TO solve a linear equation
in this site.